OpenStudy (anonymous):

Two trains leave a station at the same time, one heading east and the other west. The eastbound train travels 16 miles per hour slower than the westbound train. If the two trains are 680 miles apart after 5 hours, how many miles per hour is the eastbound train traveling?

OpenStudy (anonymous):

r*t = d Let x be the speed of the East bound train. Solve the following for x:\[((x+16)+(x))5=680\]\[x = 60 \text{ mph}\]\[76*5+60*5 = 680 \]